# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v2

@@ 4.1 The Linear Representation Hypothesis
The linear representation hypothesis suggests that human-interpretable concepts in LLMs are represented linearly. This idea is supported by empirically evidence in various forms, including the following:
Concepts as Directions: Each concept is represented as a direction determined by the differences (i.e.,, vector offset) in representations of counterfactual pairs that vary only in one latent concept of interest,, e.g., gender. For instance, Rep(‘men’)-Rep(‘women’) $\approx$ Rep(‘king’)-Rep(‘queen’) (Mikolov et al., 2013; Pennington et al., 2014; Turner et al., 2023).
Concept Manipulability: Previous studies have demonstrated that the value of a concept can be altered independently from others by introducing a corresponding steering vector (Li et al., 2024; Wang et al., 2023; Turner et al., 2023). For example, transitioning an output from a false to a truthful answer can be accomplished by adding a vector offset derived from counterfactual pairs that differ solely in the false/truthful concept. This illustrates how such vectors encapsulate the essence of false/truthful concepts.
Linear Probing: The value of a concept is often measured using a linear probe. For instance, the probability that the output language is French is logit-linear in the representation of the input. In this context, the linear weights can be interpreted as representing the concept of English/French (Park et al., 2023; Marks & Tegmark, 2023).
@@ 4.2 Understanding the Linear Representation Hypothesis
Recent works have attempted to formulate a unified framework to gain a deeper understanding of this phenomenon (Park et al., 2023; Marconato et al., 2024). However, these efforts fall short of addressing the underlying reasons for the emergence of these linear properties, largely due to a lack of connection to latent generative models. While recent research (Jiang et al., 2024) attempts to explain these linear properties through latent generative models, it has two key limitations: 1) their results rely on strong assumptions, such as the invertibility of the mapping from latent space to observed space, which, as discussed in Section 2, can be overly restrictive; and 2) their explanation is primarily framed through the lens of the implicit bias of gradient descent. In contrast, our approach does not make such strong assumptions about the mapping. Instead, we seek to explain these properties from the perspective of identifiability. Specifically, we have two main corollaries derived from our identifiability result.
@@
Assume that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$. Define the degree of approximate invertibility of the mapping from $\mathbf{c}$ to $\mathbf{x}$ by $1-p(\mathbf{c}=\mathbf{c}_{\mathbf{x}}^{*}\mid\mathbf{x})=\epsilon_{\mathbf{x}}$, where $\mathbf{c}_{\mathbf{x}}^{*}$ denotes the dominant mode of $p(\mathbf{c}\mid\mathbf{x})$. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ be two inputs that differ only in the value of the $i$th concept variable $c^{i}$, with $c^{i}=0$ for $\mathbf{x}_{0}$ and $c^{i}=1$ for $\mathbf{x}_{1}$, when $\epsilon_{\mathbf{x}}\rightarrow 0$, we have:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})-\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx s_{1}\mathbf{A}_{:,k_{1}}-s_{0}\mathbf{A}_{:,k_{0}},$
where $s_{1}\approx\log p(c^{i}=1|\mathbf{x}_{1})$, $s_{0}\approx\log p(c^{i}=0|\mathbf{x}_{0})$, and $k_{0}$ and $k_{1}$ be the indices of the concentrated mode $\mathbf{c}_{\mathbf{x}_{0}}^{*}$ for $\mathbf{x}_{0}$ and $\mathbf{c}_{\mathbf{x}_{1}}^{*}$ for $\mathbf{x}_{1}$, in the vector $\left[\log p(\mathbf{c}=\mathbf{c}_{j}\mid\mathbf{x})\right]_{\mathbf{c}_{j}}$, respectively.
@@
The corollary also supports the notion that a concept’s value can be adjusted by adding a corresponding steering vector, such as $s_{1}\mathbf{A}_{:,k_{1}}-s_{0}\mathbf{A}_{:,k_{0}}$. Adding the steering vector effectively modifies the original representation to produce a new representation, i.e., $\mathbf{\hat{f}}_{\mathbf{x}}=\mathbf{f}_{\mathbf{x}}(\mathbf{x})+\alpha(s_{1}\mathbf{A}_{:,k_{1}}-s_{0}\mathbf{A}_{:,k_{0}})=\mathbf{A}([\log p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}+\boldsymbol{\alpha}_{s})$, where $\alpha$ corresponds a weight as introduced (Wang et al., 2023; Turner et al., 2023), $\boldsymbol{\alpha}_{s}$ is a vector whose $k_{0}$-th and $k_{1}$-th elements are $-\alpha*s_{0}$ and $\alpha*s_{1}$ respectively, while all other elements are zero. Thus, manipulating a concept via a steering vector is, in essence, equivalent to modifying the posterior distribution of the concept of interest, which directly impacts the model’s output.
@@
Assume that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$. Then, for a data pair $(\mathbf{c},\mathbf{x})$, where $\mathbf{x}$ is generated by the latents $\mathbf{c}$ in the proposed latent variable model, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ is linearly separable with respect to the latents $\mathbf{c}$, by a linear classifier with weight matrix $\mathbf{W}$, where $\mathbf{W}\mathbf{A}\approx\mathbf{I}$.
The corollary supports the idea that concepts, e.g., English/French, can be predicted using a linear probe on the corresponding representations of the input. This is because the learned representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ in LLMs is a linear transformation of the log of the conditional distribution $p(\mathbf{c}|\mathbf{x})$. Such a transformation implies that a linear classifier is sufficient to align the model’s predicted distribution $p(\mathbf{c}|\mathbf{x})$ with the true distribution over classes $p(\mathbf{c})$. This alignment is achieved through the minimization of cross-entropy loss, ensuring effective prediction of the latent concepts.
@@ 4.3 Unifying the Linear Representation Hypothesis
Corollary 4.2 reveals that the weights $\mathbf{W}$ learned through linear probing are intimately connected to the linear transformation $\mathbf{A}$ from our identifiability results, with the key relationship $\mathbf{W}\mathbf{A}=\mathbf{I}$. This relationship, previously unexplored, provides a unified framework for understanding the Linear Representation Hypothesis across various manifestations. These manifestations include Concepts as Directions, Concept Manipulability, and Linear Probing, all of which are unified through the linear transformation $\mathbf{A}$. Specifically:
  • Concepts as directions and concept manipulability focus on extracting or manipulating the columns of $\mathbf{A}$, which correspond to the latent concepts of interest.
  • Linear Probing, on the other hand, involves learning the corresponding rows of $\mathbf{W}$, which can be interpreted as mapping the learned representations to the concept space.
@@
By leveraging the relationship $\mathbf{W}\mathbf{A}=\mathbf{I}$, we demonstrate that these different forms of the linear representation hypothesis are not independent, but are instead intrinsically linked through the same linear transformation $\mathbf{A}$ from our identifiability result. This unified viewpoint underscores how $\mathbf{A}$ serves as the underlying structure that connects concept as directions, concept manipulation, and linear probing.
5 Experiments on Simulation and Real Data
@@
We begin by conducting experiments on synthetic data, which is generated through the following process: First, we create random directed acyclic graphs (DAGs) with $n$ latent variables, representing concepts. For each random DAG, the conditional probabilities of each variable given its parents are modeled using Bernoulli distributions, where the parameters are sampled uniformly from $[0.2,0.8]$. To simulate a nonlinear mixture process, we then convert the latent variable samples into one-hot format and randomly apply a permutation matrix to the one-hot encoding, generating one-hot observed samples. These are then transformed into binary observed samples. To simulate next-token prediction, we randomly mask part of the binary observed data, i.e., ${x}_{i}$, and use the remaining portion, i.e., $\mathbf{x}_{j}$ where $j\neq i$, to predict the masked part ${x}_{i}$. More details can be found in Section F.1 in Appendix.
@@
Figure 2: (Left) Classification accuracy under varying numbers of observed variables: as the number increases, the latent-to-observed mapping becomes more invertible, leading to improved recovery of latent variables—consistent with our result in Theorem 3.1. (Right) Classification accuracy across different graph structures demonstrates the robustness of the identifiability results to graph structure and latent variable size.
@@
Figure 3: Results of the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ across the LLaMA-2 and Pythia model families. Here, $\mathbf{A}_{s}$ represents a matrix derived from the feature differences of 27 counterfactual pairs, while $\mathbf{W}_{s}$ is a weight matrix obtained from a linear classifier trained on these features. The product approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.2. Refer to Appendix H for more results on LLaMA-3 and DeepSeek-R1.
@@
We now present experiments on pre-trained LLMs. While it is challenging to collect all latent variables from real-world data to directly evaluate our identifiability result in Thorem 3.1, we can instead assess our corollaries to indirectly validate it. For Corollary 4.1, prior studies have already demonstrated the linear representation properties of LLM embeddings using counterfactual pairs or linear probing (Mikolov et al., 2013; Pennington et al., 2014; Turner et al., 2023; Li et al., 2024; Wang et al., 2023; Park et al., 2023). Therefore, we shift our focus to a new property in Corollary 4.2, which has not been explored in previous works. To investigate it, we utilize 27 counterfactual pairs from Park et al. (2023), which are constructed based on the Big Analogy Test dataset (Gladkova et al., 2016). More details can be found in Section F.2 in Appendix. Specifically, supported by Corollary 4.1, we compute the differences between the features of these 27 counterfactual pairs on pre-trained LLMs to construct a matrix $\mathbf{A}_{s}$ with dimensions $27\times dim$, where $dim$ corresponds to the feature dimension of the pre-trained LLM used. This matrix can be regarded as a submatrix of a variant of $\mathbf{A}$, in which each row is associated with the probability of a specific latent concept. Additionally, we train a linear classifier using the features of these counterfactual pairs and extract the corresponding weights, supported by Corollary 4.2, forming a weight matrix $\mathbf{W}_{s}$ with dimensions $dim\times 27$, which also serves as a submatrix of $\mathbf{W}$. Finally, we examine the product of $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$ to assess whether it approximates the identity matrix. Figure 3 displays the results of this product across the LLaMA-2 and Pythia model families. The results show that the product approximates the identity matrix, which is consistent with the theoretical finding in Corollary 4.2.
6 Conclusion and Discussion

### v3

@@ 4.1 The Linear Representation Hypothesis
The linear representation hypothesis suggests that human-interpretable concepts in LLMs are represented linearly. This idea is supported by empirically evidence in various forms, including:
Concepts as Directions: Each concept is represented as a direction determined by the differences (i.e.,, vector offset) in representations of pairs that vary only in one latent concept of interest,, e.g., gender. For instance, Rep(‘men’)-Rep(‘women’) $\approx$ Rep(‘king’)-Rep(‘queen’) [54, 60, 72].
Concept Manipulability: Previous studies have demonstrated that the value of a concept can be altered independently from others by introducing a corresponding steering vector [43, 76, 72]. For example, transitioning an output from a false to a truthful answer can be accomplished by adding a vector offset derived from counterfactual pairs that differ solely in the false/truthful concept. This illustrates how such vectors encapsulate the essence of false/truthful concepts.
Linear Probing: The value of a concept is often measured using a linear probe. For instance, the probability that the output language is French is logit-linear in the representation of the input. In this context, the linear weights can be interpreted as representing the concept of English/French [59, 52].
@@ 4.2 Understanding the Linear Representation Hypothesis
The linear representation hypothesis has received growing empirical support in recent years. While recent work has aimed to develop unified frameworks for a deeper understanding of this phenomenon [59, 51, 35], our approach seeks to explain it through the lens of identifiability. Before that, we first provide the following definition:
Definition 4.1.
We define the degree of approximate invertibility of the mapping from $\mathbf{c}$ to $\mathbf{x}$ by introducing an error term $\epsilon_{\mathbf{x}}$, as follows: $1-p(\mathbf{c}=\mathbf{c}^{*}|\mathbf{x})=\epsilon_{\mathbf{x}}$, where $10<\epsilon_{\mathbf{x}}<$ 1, and $\mathbf{c}^{*}=\arg\max_{\mathbf{c}}p(\mathbf{c}|\mathbf{x})$ represents the dominant mode of $p(\mathbf{c}|\mathbf{x})$.
We then present two key corollaries derived from our identifiability results.
@@
Suppose that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}+\mathbf{b}$. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ be a pair of inputs that differ only in the $i$-th concept variable $c^{i}$. Then when $\epsilon_{\mathbf{x}}\to 0$ in Def. 4.1, $\mathbf{f}_{x}(\mathbf{x}_{1})-\mathbf{f}_{x}(\mathbf{x}_{0})\approx\tilde{\mathbf{A}}^{i}\left(\left[\log p(c^{i}\mid\mathbf{x}_{1})-\log p(c^{i}\mid\mathbf{x}_{0})\right]_{c^{i}}\right)$, where $\tilde{\mathbf{A}}^{i}=\mathbf{A}\mathbf{B}^{i}$, where $\mathbf{B}^{i}$ is a binary lifting matrix that broadcasts each entry of $[\log p(c^{i}\mid\mathbf{x})]_{c^{i}}$ to the corresponding index in $[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})]_{\mathbf{c}_{i}}$.
@@
[…]
@@
Suppose that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}+\mathbf{b}$. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ be pair data that differ only in the $i$-th concept variable $c^{i}$, with labels $c^{i}$. Then when $\epsilon_{\mathbf{x}}\to 0$ in Def. 4.1, the corresponding representations $(\mathbf{f}(\mathbf{x}_{0}),\mathbf{f}(\mathbf{x}_{1}))$ are linearly separable with a weight matrix $\mathbf{W}$ satisfying $\mathbf{W}\tilde{\mathbf{A}}^{i}\approx s\mathbf{I}$, where $s$ accounts for softmax scaling. The corresponding logit is the unnormalized $[p({c}^{i}\mid\mathbf{x})]_{{c}^{i}}$.
This corollary supports that latent concepts, e.g., English vs. French, can be reliably predicted using a linear probe on the model’s internal representations. This linear separability enables alignment between the model’s predictive distribution and and the true class distribution, achieved via cross-entropy minimization. A specific analysis for the binary case of concept $c^{i}$ is provide in Appendix F.2.
@@ 4.3 Unifying the Linear Representation Hypothesis
Corollary 4.3 reveals that the weights $\mathbf{W}$ learned through linear probing are intimately connected to the linear transformation $\mathbf{A}$ from our identifiability results, with the key relationship $\mathbf{W}\tilde{\mathbf{A}}^{i}\approx s\mathbf{I}$. This relationship, previously unexplored, provides a unified framework for understanding the linear representation hypothesis across various forms, e.g., concepts as directions, concept manipulability, and linear probing. By leveraging the relationship $\mathbf{W}\tilde{\mathbf{A}}^{i}\approx s\mathbf{I}$, we demonstrate that these different forms of the linear representation hypothesis are not independent, but are instead intrinsically linked through the same linear transformation $\mathbf{A}$ from our identifiability result. This unified viewpoint underscores how $\mathbf{A}$ serves as the underlying structure that connects concept as directions, concept manipulation, and linear probing.
5 Evaluating Sparse Autoencoders and Structured Sparse Autoencoders
Evaluating SAEs
Broadly speaking, SAEs are designed with two primary objectives. First, they aim to learn a set of latent features $\mathbf{z}$ such that sparse linear combinations $\boldsymbol{\beta}\mathbf{z}$ can accurately reconstruct internal representations from LLMs, i.e., $\mathbf{f}(\mathbf{x})\approx\boldsymbol{\beta}\mathbf{z}.$ Second, they seek to ensure that each learned feature $z_{i}$ corresponds to a disentangled, human-interpretable concept, thereby enabling a mechanistic understanding of LLMs. While reconstruction loss is commonly used to assess how well representations are reconstructed [61, 62, 27, 11], it is a limited proxy for the second objective. A key challenge lies in the absence of ground truth for the underlying concepts [37], making the evaluation of feature disentanglement particularly difficult.
To address it, we propose a new evaluation method for SAEs grounded in our theoretical insights. Specifically, based on Theorem 3.1, the LLM representations $\mathbf{f}_{x}(\mathbf{x})$ can be approximated by $\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}$. Meanwhile, SAEs are trained to reconstruct the same $\mathbf{f}_{x}(\mathbf{x})$ by $\boldsymbol{\beta}\mathbf{z}$. Combining the two, we arrive at: $\boldsymbol{\beta}\mathbf{z}\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}.$ This suggests that SAE features $\mathbf{z}$ are linearly related to $\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}$. Based on this, if we expect that each $z_{i}$ encodes a single concept $c^{i}$, then it naturally follows that $z_{i}$ should be linearly related only to the posterior of a single concept $\log p(c^{i}\mid\mathbf{x})$. Consequently, we can evaluate whether each $z_{i}$ has successfully learned a disentangled concept by measuring its linear correlation with $\log p(c^{i}\mid\mathbf{x})$. The question is how can we obtain $p(c^{i}\mid\mathbf{x})$.
Based on Corollary 4.3, we can obtain $p(c^{i}\mid\mathbf{x})$ by pair data that differ only in the $i$-th concept variable $c^{i}$. Specifically, we can construct paired data $(\mathbf{x}_{0},\mathbf{x}_{1})$ that differ in only a single binary concept $c^{i}$, with labels $c^{i}=0$ for $\mathbf{x}_{0}$ and $c^{i}=1$ for $\mathbf{x}_{1}$. We then train a linear classifier (i.e., linear probing) in a supervised manner on the corresponding LLM representations $\mathbf{f}(\mathbf{x}_{0})$ and $\mathbf{f}(\mathbf{x}_{1})$, with the goal of predicting $c^{i}$. Once trained, the corresponding logit provides a reliable estimate of the posterior probability $p(c^{i}=1\mid\mathbf{x})$. See Appendix F.2 for a binary special case of Corollary 4.3.
Structured SAEs
Again, building on $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}$ in Theorem 3.1, SAEs attempt to recovering interpretable concepts from entangled representations in a unsupervised manner, which is typically an ill-posed inverse problem. As a result, using sparsity-inducing prior only, e.g., $\ell_{p}$ [38] or top-$k$ regularization [26, 17], may be insufficient for regularizing the solution space. To address this limitation, by recognising concepts may appear dependencies, we propose structured SAEs, which incorporate additional constraints to better reflect the inductive biases of the latent concept space by accounting for potential dependencies among concepts. In particular, we introduce low-rank regularization to complement sparsity-based priors (Other forms of structured regularization may also be worth exploring.). Formally, structured SAEs minimize the following objective:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{train}}}\left[\|\mathbf{f}(\mathbf{x})-\mathbf{\bar{f}}(\mathbf{x})\|_{2}^{2}+\lambda_{t}\left(\|\mathbf{S}\|_{p_{t}}^{p_{t}}+\gamma\|\mathbf{R}\|_{\text{nuc}}\right)\right],$ (6)
where $\mathbf{\bar{f}}(\mathbf{x})$ denotes reconstruction, $\lambda_{t}$ is the dynamically adjusted sparsity coefficient at step $t$, following the $p$-annealing SAE strategy [38], $\gamma$ is a hyperparameter that balances the sparsity penalty and the low-rank regularization, $\|\mathbf{S}\|_{p_{t}}^{p_{t}}=\sum_{i}|S_{i}|^{p_{t}}$ is the adaptive $\ell_{p_{t}}$ norm promoting sparsity, $\|\mathbf{R}\|_{\text{nuc}}$ is the nuclear norm, used to encourage low-rank structure on $\mathbf{R}$, and $\mathbf{z}=\mathbf{S}+\mathbf{R}$.
6 Empirical Evaluation on Simulated and Real Data with LLMs
@@
We begin by conducting experiments on synthetic data, which is generated through the following process: First, we create random directed acyclic graphs (DAGs) with $n$ latent variables, representing concepts. For each random DAG, the conditional probabilities of each variable given its parents are modeled using Bernoulli distributions, where the parameters are sampled uniformly from $[0.2,0.8]$. To simulate a nonlinear mixture process, we then convert the latent variable samples into one-hot format and randomly apply a permutation matrix to the one-hot encoding, generating one-hot observed samples. These are then transformed into binary observed samples. To simulate next-token prediction, we randomly mask part of the binary observed data, i.e., ${x}_{i}$, and use the remaining portion, i.e., $\mathbf{x}_{j}$ where $j\neq i$, to predict the masked part ${x}_{i}$. Refer to Section G.1 in Appendix for more details.
Figure 2: (Left) Classification accuracy under varying numbers of observed variables. (Right) Classification accuracy across different graph structures.
@@
Figure 3: Results of the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ across the LLaMA-2 and Pythia model families. Here, $\mathbf{A}_{s}$ represents a matrix derived from the feature differences of 27 counterfactual pairs, while $\mathbf{W}_{s}$ is a weight matrix obtained from a linear classifier trained on these features. The product approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.3.
@@
We now present experiments on pre-trained LLMs. While it is challenging to collect all latent variables from real-world data to directly evaluate our identifiability result in Thorem 3.1, we can instead assess our corollaries to indirectly validate it. For Corollary 4.2, prior studies have already demonstrated the linear representation properties of LLM embeddings using counterfactual pairs that differ only in a single concept of interest [54, 60, 72, 43, 76, 59]. Therefore, we shift our focus to a new property highlighted in Corollary 4.3, specifically the relationship $\mathbf{W}\tilde{\mathbf{A}}^{i}\approx s\mathbf{I}$, which has not been explored in prior work. To investigate it, we utilize 27 counterfactual pairs from [59] that differ only in a binary concept, which are constructed based on the Big Analogy Test dataset [28]. More details can be found in Section G.2 in Appendix. Specifically, supported by Corollary 4.2, we compute the differences between the representations of these 27 counterfactual pairs on pre-trained LLMs to construct a matrix $\mathbf{A}_{s}$ with dimensions $27\times dim$, where $dim$ corresponds to the representation dimension of the pre-trained LLM used. $\mathbf{A}_{s}$ denote a variant of $\tilde{\mathbf{A}}^{i}$, with each row is associated with the probability of a specific binary concept. See a binary special case of Corollary 4.2 in Appendix F for more details. Additionally, we train a linear classifier using the representations of these counterfactual pairs and extract the corresponding weights, supported by Corollary 4.3, forming a weight matrix $\mathbf{W}_{s}$ with dimensions $dim\times 27$. Finally, after normalizing $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$ to avoid scaling $s$ in Corollary 4.3, we examine the product of $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$ to assess whether $\mathbf{A}_{s}\mathbf{W}_{s}\approx\mathbf{I}$. Figure 3 displays the results of this product across the LLaMA-2 and Pythia model families. Refer to Appendix J for more results on LLaMA-3 and DeepSeek-R1. The results show that the product approximates the identity matrix, which is consistent with the theoretical finding in Corollary 4.3.
Experiments on Structured SAEs
We train four sparse SAEs variants—top-$k$ SAE [26], batch‑top‑$k$ SAE [17], $p$-annealing SAE [38], and the proposed structured SAE. Following Theorem 3.1, each SAE is trained on representations from the final hidden layer of the 70 million-parameter Pythia language model [8], with The Pile corpus [25]. For evaluation, we use 27 counterfactual pairs from [59] again to train a linear classification using LogisticRegression classifier from the scikit-learn library, to obtain logits, i.e., unnormalized $p(c^{i}=1|\mathbf{x})$, for each of the 27 pairs. These counterfactual pairs are also passed through the trained SAEs to extract features $\mathbf{z}$. We search for the best-matching feature $z_{i}$ for each $p(c^{i}|\mathbf{x})$ according to Pearson correlation between $\exp(z_{i})$ and $p(c^{i}|\mathbf{x})$. Table 4(a) shows the Pearson correlations for a subset of 27 concepts, with additional results provided in Appendix H. These findings highlight the advantage of the proposed structured SAE, which benefits from incorporating structured regularization. This advantage is further evident in the reconstruction loss, as illustrated in Figure 4(b).
Pearson Correlation ($\uparrow$)
Concepts top-$k$ batch-top-$k$ $p$-annealing Ours
3psg_ved 0.432 0.332 0.680 0.691
adj_adj_+_ly 0.379 0.396 0.794 0.833
adj_comparative 0.425 0.334 0.592 0.727
average 0.365 0.331 0.638 0.674
country_capital 0.205 0.165 0.495 0.540
french_german 0.264 0.226 0.550 0.607
french_spanish 0.389 0.264 0.560 0.561
frequent_infrequent 0.262 0.255 0.472 0.554
(a) Pearson correlations across 4 SAEs.
(b) Reconstruction loss during training
Figure 4: Comparison of SAE models: correlation scores and reconstruction loss on the validation dataset.
7 Conclusion